# Data-Driven Superstabilizing Control under Quadratically-Bounded Errors-in-Variables Noise

### v1

@@ Abstract
The Error-in-Variables model of system identification/control involves nontrivial input and measurement corruption of observed data, resulting in generically nonconvex optimization problems. This paper performs full-state-feedback stabilizing control of all discrete-time linear systems that are consistent with observed data for which the input and measurement noise obey quadratic bounds. Instances of such quadratic bounds include elementwise norm bounds (at each time sample), energy bounds (across the entire signal), and chance constraints arising from (sub)gaussian noise. Superstabilizing controllers are generated through the solution of a sum-of-squares hierarchy of semidefinite programs. A theorem of alternatives is employed to eliminate the input and measurement noise process, thus improving tractability. Effectiveness of the scheme is generated on an example system in the chance-constrained set-membership setting where the input and state-measurement noise are i.i.d. normally distributed.
@@ I Introduction
This paper proposes a method to stabilize linear systems corrupted by quadratically bounded Errors-in-Variables (EIV) noise [1]. State and input observations $\mathcal{D}=\{\hat{x}_{t},\hat{u}_{t}\}_{t=0}^{T}$ are collected along a trajectory of an $n$-state $m$-input linear system
$x_{t+1}$ $=Ax_{t}+Bu_{t}$ (1a)
@@
in which the collected data in $\mathcal{D}$ is corrupted by state noise $\Delta x\in\mathbb{R}^{nT}$ and input noise $\Delta u\in\mathbb{R}^{m(T-1)}$. This paper focuses on the setting where the noise processes $(\Delta x,\Delta u)$ satisfy a collection of $N$ convex quadratic constraints as
@@
in which the known constraint matrices $F_{j}$ and $G_{j}$ have compatible dimensions. These quadratic constraints could arise from deterministic knowledge of $(\Delta x,\Delta u)$, or from high-probability chance-constraints imposed on stochastic processes $(\Delta x,\Delta u)$ if a robust description is overly conservative [2, 3]. Our goal is to find a gain matrix $K$ such that the full-state-feedback control policy $u_{t}=Kx_{t}$ can simultaneously stabilize all plants $(A,B)$ that are consistent with the data in $\mathcal{D}$ under the noise description in (2).
@@
This paper follows the framework of set-membership direct Data Driven Control (DDC). In direct DDC, a control policy is formed from the collected data and modeling assumptions without first performing system identification (and synthesizing a controller for the identified system) [4]. Set-membership DDC has three main ingredients: the set of data-consistent plants (given a noise model), the set of commonly stabilized plants by a designed controller, and the certificate of set-containment that the stabilized-set contains the consistent-set [5]. The Matrix S-Lemma can be used to provide proofs of quadratic and worst-case $H_{2}$ or $H_{\infty}$ robust control when the noise model is defined by a matrix ellipsoid (quadratic matrix inequality) [6, 7]. Farkas-based certificates for polytope-in-polytope membership have been used for robust superstabilization [8] and positive-stabilization [9]. Sum of Squares (SOS) certificates of nonnegativity have been employed for stabilization of more general nonlinear systems [10, 11]. We note that other non-set-membership DDC methods include using Virtual Reference Feedback Tuning [12] and Willem’s Fundamental Lemma [13].
@@
Most DDC methods focus solely on process-noise corruption, allowing for the synthesis of controllers through the solution of computationally simple convex programs. This paper continues a line of work in addressing the more challenging setting of EIV superstabilization, proposing a method that can handle the more difficult but realistic EIV noise scenario at the cost of more expensive computational requirements. Our prior work on superstabilization [14, 15] of EIV systems includes full-state-feedback for polytope-bounded noise [16] (e.g. $L_{\infty}$-bounds) and dynamic output feedback for SISO plants [17]. In this work, we will ensure superstabilization under quadratically bounded noise. This will involve developing matrix-SOS expressions defined for (multiple) quadratic constraints in (2). Computational complexity is reduced by eliminating $(\Delta x,\Delta u)$ using a Theorem of Alternatives [18]. The concurrent and independently developed similar work in [19] performs lossless quadratic stabilization in the presence of a single quadratic-matrix-inequality-representable quadratic constraint, with conservatism added under multiple quadratic constraints.
@@
This paper has the following structure: Section II reviews preliminaries including notation, quadratically constrained noise, superstabilization, and matrix-SOS methods. Section III presents infinite-dimensional Linear Programs to perform superstabilization under quadratically-bounded EIV noise, and applies a Theorem of Alternative to eliminate the noise variables $(\Delta x,\Delta u)$ from the linear inequalities. Section IV truncates the infinite-dimensional LPs using the moment-SOS hierarchy of Semidefinite Programs, and tabulates computational complexity of different approaches. Section V incorporates recorded process noise into the quadratically-bounded framework. Section VI demonstrates our method for EIV-tolerant superstabilizing control of example systems. Section VII concludes the paper.
@@ II-A Notation
$a..b$
natural numbers between $a$ and $b$ (inclusive)
$\mathbb{R}_{\geq 0}$
nonnegative orthant
$\pi^{x}$
projection operator $(x,y)\rightarrow x$
$A^{\top}$
Transpose of matrix $A$
$A^{+}$
Pseudoinverse of matrix $A$
$\mathbb{L}^{n}$
$n$-dimensional Second-Order Cone (SOC)
@@
$\mathcal{N}(\mathbf{m},\Sigma)$
Normal distribution with mean $\mathbf{m}$
@@
$\mathbb{R}[x]$
polynomials in indeterminate $x$
$\deg(h)$
degree of polynomial $h\in\mathbb{R}[x]$
$\mathbb{S}^{r}[x]$
$r\times r$ polynomial-valued-matrices in $x$
$\Sigma^{r}[x]$
SOS polynomial-valued-matrices in $x$
$\langle A,B\rangle$
Matrix pairing $\textrm{Tr}(A^{\top}B)=\sum_{ij}A_{ij}B_{ij}$
@@ II-B1 Elementwise Norm Constraints
is $P_{\textrm{joint}}=(1-\delta_{x})^{T}(1-\delta_{u})^{T-1}$. A controller $u=Kx$ certifiably stabilizes all plants consistent with (I) and (II-B1) will be able to stabilize the true system with probability $P_{\textrm{joint}}$. Similar elementwise quadratic chance-constraints arise when $\Delta x,\Delta u$ are drawn from i.i.d. subgaussian distributions [20].
@@ II-B2 Energy Constraints
is compatible with the structure of (2) if $(Q,R,Q_{T})$ are all Positive Semidefinite (PSD). The mixed structure (for appropriately dimensioned $\tilde{Q},\tilde{R},\tilde{S}$)
$J_{\textrm{mix}}=\begin{bmatrix}\Delta x\\ \Delta u\end{bmatrix}^{T}\begin{bmatrix}\tilde{Q}&\tilde{S}\\ \tilde{S}^{\top}&\tilde{R}\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta u\end{bmatrix}$ (6)
can also be posed as the left-hand-side of the (2) inequality, thus allowing for energy constraints on the noise $(\Delta x,\Delta u)$.
@@ II-C Superstabilizing Control
$\lVert W^{+}AW\rVert_{\infty}<1\quad\text{($L_{\infty}$ Operator Norm).}$). (7)
The system is superstable if it is $W$-superstable with $W=I_{n}$. Any superstable system $x_{t+1}=Ax_{t}$ obeys the the decay bound of $\lVert x_{t}\rVert_{\infty}\leq\lVert A\rVert_{\infty}^{t}\lVert x_{0}\rVert_{\infty}$. This decay bound generalizes to $W$-superstability as in $\lVert Wx_{t}\rVert_{\infty}\leq\lVert W^{+}AW\rVert_{\infty}^{t}\lVert Wx_{0}\rVert_{\infty}$. $W$-superstabilization of the system $x_{t+1}=Ax_{t}+Bu_{t}$ proceeds by choosing a gain $K\in\mathbb{R}^{m\times n}$ with $u_{t}=Kx_{t}$ such that $\lVert A+BK\rVert_{\infty}<1$. The superstabilization problem of minimizing the decay bound is a finite-dimensional LP:
@@
$\textstyle\lvert[W^{+}(A+BK)W]_{ij}\rvert\leq M_{ij}$ $\forall i,j\in 1..f.$ (8b)
@@ II-D Sum-of-Squares Matrices Background
The matrix $p\in\mathbb{S}^{r}[x]$ satisfies \@iaciPMI Polynomial Matrix Inequality (PMI) over the region in (10) if $\forall x\in\mathbb{K}:\ p(x)\succeq 0.$ A sufficient condition for this PMI to hold is if there exists SOS-matrices $\{\sigma_{j}(x)\}_{j=0}^{N_{c}}$ such that
$p(x)$ $=\sigma_{0}(x)+\textstyle\sum_{i=j}^{N_{c}}\langle g_{j}(x),\sigma_{j}(x)\rangle$ (11a)
@@
The set of matrices in $\mathbb{S}^{r}[x]$ possessing a representation as in (11a) (existence of $\{\sigma_{j}\}$) will be written as the Weighted Sum of Squares (WSOS) set $\Sigma^{r}[\mathbb{K}]$. The degree-$2k$-bounded WSOS cone $\Sigma^{r}[\mathbb{K}]_{\leq 2k}$ is the set of matrices with representation in (II-D) such that $\deg\sigma_{0}\leq 2k$ and $\forall j\in 1..N_{c}:\ \deg\langle g_{j}(x),\sigma_{j}(x)\rangle\leq 2k$. The representation of $\mathbb{K}$ by polynomial matrices in (10) is Archimedean if there exists an $R>0$ such that the scalar polynomial $p_{R}(x)=R-\lVert x\rVert_{2}^{2}$ satisfies $p_{R}\in\Sigma^{r}[\mathbb{K}]$. If the representation in (10) is Archimedean, then every $p\in\mathbb{S}^{r}[x]$ such that $p(x)$ is Positive Definite over $K$ satisfies $(p(x)-\epsilon I_{r})\in\Sigma^{r}[K]$ for some $\epsilon>0$ (Corollary 1 of [22]). Testing membership of $p\in\Sigma^{r}[x]_{\leq 2k}$ through (II-D) can be accomplished by solving \@iaciLMI Linear Matrix Inequality (LMI) in the Gram matrices for $\{\sigma_{j}\}$ under coefficient-matching equality constraints. A common choice of polynomial vector $v(x)$ used to represent the SOS matrices in (9) is the vector of all $\binom{n+k}{k}$ monomials of degree $\leq k$. The maximal-size PSD constraint involved in (II-D) under the monomial choice for $v(x)$ is $r\binom{n+k}{k}$.
@@ III-A Consistency Set
Define $h^{0}$ as the following residual:
$h_{t}^{0}$ $=\hat{x}_{t+1}-A\hat{x}_{t}-B\hat{u}_{t}$ $\forall t\in 1..T-1.$ (12)
The quantity $h^{0}$ will be equal to zero at all $t\in 1..T-1$ if there is no noise present in the system.
The joint consistency set $\bar{\mathcal{P}}(A,B,\Delta x,\Delta u)\subset\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times m}\times\mathbb{R}^{nT}\times\mathbb{R}^{m(T-1)}$ of plants and noise values consistent with the dynamics (1a) and noise properties (2) is
@@
The formulation (15) involves bilinear terms $A\Delta x$ and $B\Delta u$, which could lead to nonconvex or possibly disconnected sets $\bar{\mathcal{P}}$ [23]. By defining a matrix $\Xi_{A}\in\mathbb{R}^{n(T-1)\times nT}$ with
$\Xi$ $=\begin{bmatrix}A&-I&\mathbf{0}&\ldots\\ \mathbf{0}&A&-I&\ldots\\ \vdots&\vdots&\vdots&\ddots\end{bmatrix}$ (19)
$=[(I_{T-1}\otimes A),\mathbf{0}_{n(T-1)\times n}]+[\mathbf{0}_{n(T-1)\times n},-I_{n(T-1)}],$
and defining $n_{\ell}$ to be the column dimension of $F_{\ell}$ and $G_{\ell}$ for each $\ell$ ($F_{\ell}\in\mathbb{R}^{n_{\ell}\times nT},\ G_{\ell}\in\mathbb{R}^{n_{\ell}\times m(T-1)}$, the consistency set in (15) can be equivalently expressed as
$\bar{\mathcal{P}}:=\ \begin{Bmatrix}[l]\Xi\Delta x+(I_{T-1}\otimes B)\Delta u+h_{0}=0\\ (F_{\ell}\Delta x+G_{\ell}\Delta u,1)\in\mathbb{L}^{n_{\ell}}&\forall\ell\in 1..L\\ \end{Bmatrix}.$ (22)
@@ III-B Full Linear Program
$\operatorname*{find}_{M,K}$ $\forall(A,B,\Delta x,\Delta u)\in\bar{\mathcal{P}}(A,B,\Delta x,\Delta u):$ (24a)
@@
$\qquad M_{ij}(A,B)-(W^{+}\left(A+BK\right)W)_{ij}\geq 0$
$\qquad M_{ij}(A,B)+(W^{+}\left(A+BK\right)W)_{ij}\geq 0$
@@
The $W$-superstabilizing linear system in (II-C) is fulfilled for each $(A,B)\in\mathcal{P}$, as expressed by the $\forall(A,B,\Delta x,\Delta u)\in\bar{\mathcal{P}}$ quantification in (24a). ∎
Theorem III.2
Under the assumption that $\bar{\mathcal{P}}$ is compact, the function $M$ can be selected to be continuous (Lemma 3.1 of [24]) and even polynomial (Lemma 3.2 of [24])
Proof:
This follows from Lemmas 3.1 (continuity) and 3.2 (polynomial approximability) of [24]. The same type of linear system structure for superstabilization in (24b)-(24d) is used (with $W=I_{n}$), but the set $\bar{\mathcal{P}}$ in [24] is defined by polytopic-bounded noise rather than quadratically-bounded noise (15). Proof details are omitted for space. ∎

### v2

@@ Abstract
The Error-in-Variables model of system identification/control involves nontrivial input and measurement corruption of observed data, resulting in generically nonconvex optimization problems. This paper performs full-state-feedback stabilizing control of all discrete-time linear systems that are consistent with observed data for which the input and measurement noise obey quadratic bounds. Instances of such quadratic bounds include elementwise norm bounds (at each time sample), energy bounds (across the entire signal), and chance constraints arising from (sub)gaussian noise. Superstabilizing controllers are generated through the solution of a sum-of-squares hierarchy of semidefinite programs. A theorem of alternatives is employed to eliminate the input and measurement noise process, thus improving tractability.
@@ I Introduction
This paper proposes a method to stabilize linear systems corrupted by quadratically bounded Errors-in-Variables (EIV) noise [1]. State and input observations $\mathcal{D}=\{\hat{x}_{t},\hat{u}_{t}\}_{t=1}^{T}$ are collected along a trajectory of an a-priori unknown $n$-state $m$-input linear system $(A_{\star},B_{\star})$:
$x_{t+1}$ $=A_{\star}x_{t}+B_{\star}u_{t}$ (1a)
@@
in which the collected data in $\mathcal{D}$ is corrupted by state noise $\Delta x\in\mathbb{R}^{nT}$ and input noise $\Delta u\in\mathbb{R}^{m(T-1)}$. This paper focuses on the setting where the noise processes $(\Delta x,\Delta u)$ satisfy a collection of $L$ convex quadratic constraints as
@@
in which the known constraint matrices $F_{j}$ and $G_{j}$ have compatible dimensions. These quadratic constraints could arise from deterministic knowledge of $(\Delta x,\Delta u)$, or from high-probability chance-constraints imposed on stochastic processes $(\Delta x,\Delta u)$ if a robust description is overly conservative [2, 3]. Our goal is to find a gain matrix $K$ such that the full-state-feedback control policy $u_{t}=Kx_{t}$ can simultaneously stabilize all plants that are consistent with the data in $\mathcal{D}$ under the noise description in (2).
@@
This paper follows the framework of set-membership direct Data Driven Control (DDC). In direct DDC, a control policy is formed from the collected data and modeling assumptions without first performing system identification (and synthesizing a controller for the identified system) [4]. Set-membership DDC has three main ingredients: the set of data-consistent plants (given a noise model), the set of commonly stabilized plants by a designed controller, and the certificate of set-containment that the stabilized-set contains the consistent-set [5]. The Matrix S-Lemma can be used to provide proofs of quadratic and worst-case $H_{2}$ or $H_{\infty}$ robust control when the noise model is defined by a matrix ellipsoid (quadratic matrix inequality) [6, 7]. Farkas-based certificates for polytope-in-polytope membership have been used for robust superstabilization [8] and positive-stabilization [9]. Sum-of-Squares (SOS) certificates of nonnegativity have been employed for stabilization of more general nonlinear systems [10, 11]. We note that other non-set-membership DDC methods include using Virtual Reference Feedback Tuning [12] and Willem’s Fundamental Lemma [13].
@@
Most DDC methods focus solely on process-noise corruption, allowing for the synthesis of controllers through the solution of computationally simple convex programs. This paper continues a line of work in addressing the more challenging setting of EIV superstabilization, proposing a method that can handle the more difficult but realistic EIV noise scenario at the cost of more expensive computational requirements. Prior work on superstabilization [14, 15] of EIV systems includes full-state-feedback for polytope-bounded noise [16] (e.g. $L_{\infty}$-bounds) and dynamic output feedback for SISO plants [17]. In this work, we will ensure superstabilization under quadratically bounded noise. This will involve developing matrix-SOS expressions defined for (multiple) quadratic constraints in (2). Computational complexity is reduced by eliminating $(\Delta x,\Delta u)$ using a Theorem of Alternatives [18]. The concurrent and independently developed similar work in [19] performs lossless quadratic stabilization in the presence of a single quadratic-matrix-inequality-representable quadratic constraint, with conservatism added under multiple quadratic constraints.
@@
This letter has the following structure: Section II reviews preliminaries including notation, quadratically constrained noise, superstabilization, and matrix-SOS methods. Section III presents infinite-dimensional Linear Programs to perform superstabilization under quadratically-bounded EIV noise, and applies a Theorem of Alternative to eliminate the noise variables $(\Delta x,\Delta u)$ from the linear inequalities. Section IV truncates the infinite-dimensional LPs using the moment-SOS hierarchy of Semidefinite Programs, and tabulates computational complexity of different approaches. Section V describes extended superstabilization [20] of data-consistent systems. Section VI demonstrates our method for EIV-tolerant superstabilizing control of example systems. Section VII concludes the letter.
@@ II-A Notation
$a..b$ natural numbers between $a$ and $b$ (inclusive)
$\mathbb{R}_{\geq 0}\ (\mathbb{R}_{>0})$ nonnegative (positive) orthant
$\pi^{x}$ projection operator $(x,y)\rightarrow x$
$A^{\top}$ Transpose of matrix $A$
$A^{+}$ Pseudoinverse of matrix $A$
$\mathbb{L}^{n}$ $n$-dimensional Second-Order Cone (SOC)
@@
$\mathcal{N}(\mathbf{m},\Sigma)$ Normal distribution with mean $\mathbf{m}$
@@
$\mathbb{R}[x]$ polynomials in indeterminate $x$
$\deg(h)$ degree of polynomial $h\in\mathbb{R}[x]$
$\mathbb{S}^{r}[x]$ $r\times r$ polynomial-valued-matrices in $x$
$\Sigma^{r}[x]$ SOS polynomial-valued-matrices in $x$
$\langle A,B\rangle$ Matrix pairing $\textrm{Tr}(A^{\top}B)=\sum_{ij}A_{ij}B_{ij}$
@@ II-B1 Elementwise Norm Constraints
is $P_{\textrm{joint}}=(1-\delta_{x})^{T}(1-\delta_{u})^{T-1}$. A controller $u=Kx$ certifiably stabilizes all plants consistent with (1) and (4) will be able to stabilize the true system with probability $P_{\textrm{joint}}$. Similar elementwise quadratic chance-constraints arise when $\Delta x,\Delta u$ are drawn from i.i.d. subgaussian distributions [21].
@@ II-B2 Energy Constraints
is compatible with the structure of (2) if the cost matrices $(Q,R,Q_{T})$ are all Positive Semidefinite (PSD).
@@ II-C Superstabilizing Control
$\lVert WAW^{+}\rVert_{\infty}<1\quad\text{($L_{\infty}$ Operator Norm).}$). (6)
The system is superstable if it is $W$-superstable with $W=I_{n}$. Any superstable system $x_{t+1}=Ax_{t}$ obeys the the decay bound of $\lVert x_{t}\rVert_{\infty}\leq\lVert A\rVert_{\infty}^{t}\lVert x_{0}\rVert_{\infty}$. This decay bound generalizes to $W$-superstability as in $\lVert Wx_{t}\rVert_{\infty}\leq\lVert WAW^{+}\rVert_{\infty}^{t}\lVert Wx_{0}\rVert_{\infty}$. $W$-superstabilization of the system $x_{t+1}=Ax_{t}+Bu_{t}$ proceeds by choosing a gain $K\in\mathbb{R}^{m\times n}$ with $u_{t}=Kx_{t}$ such that $\lVert W(A+BK)W^{+}\rVert_{\infty}<1$. The $W$-superstabilization problem of minimizing the decay bound (for fixed $W$) is a finite-dimensional LP:
@@
$\textstyle\lvert[W(A+BK)W^{+}]_{ij}\rvert\leq M_{ij}$ $\forall i,j\in 1..f.$ (7b)
@@ II-D Sum-of-Squares Matrices Background
The matrix $p\in\mathbb{S}^{r}[x]$ satisfies \@iaciPMI Polynomial Matrix Inequality (PMI) over the region in (9) if $\forall x\in\mathbb{K}:\ p(x)\succeq 0.$ A sufficient condition for this PMI to hold is that there exist SOS-matrices $\{\sigma_{j}(x)\}_{j=0}^{N_{c}}$ and an $\epsilon>0$ such that
$p(x)$ $=\sigma_{0}(x)+\textstyle\sum_{i=j}^{N_{c}}\langle g_{j}(x),\sigma_{j}(x)\rangle+\epsilon I$ (10a)
@@
The set of matrices in $\mathbb{S}^{r}[x]$ possessing a representation as in (10a) (existence of $\{\sigma_{j}\}$) will be written as the Weighted Sum of Squares (WSOS) set $\Sigma^{r}[\mathbb{K}]$. The degree-$2k$-bounded WSOS cone $\Sigma^{r}[\mathbb{K}]_{\leq 2k}$ is the set of matrices with representation in (10) such that $\deg\sigma_{0}\leq 2k$ and $\forall j\in 1..N_{c}:\ \deg\langle g_{j}(x),\sigma_{j}(x)\rangle\leq 2k$. The representation of $\mathbb{K}$ by polynomial matrices in (9) is Archimedean if there exists an $R>0$ such that the scalar polynomial $p_{R}(x)=R-\lVert x\rVert_{2}^{2}$ satisfies $p_{R}\in\Sigma^{r}[\mathbb{K}]$. If the representation in (9) is Archimedean, then every $p\in\mathbb{S}^{r}[x]$ such that $p(x)$ is Positive Definite over $K$ satisfies $(p(x)-\epsilon I_{r})\in\Sigma^{r}[K]$ for some $\epsilon>0$ (Corollary 1 of [23]). Testing membership of $p\in\Sigma^{r}[x]_{\leq 2k}$ through (10) can be accomplished by solving \@iaciLMI Linear Matrix Inequality (LMI) in the Gram matrices for $\{\sigma_{j}\}$ under coefficient-matching equality constraints. A common choice of polynomial vector $v(x)$ used to represent the SOS matrices in (8) is the vector of all $\binom{n+k}{k}$ monomials of degree $\leq k$. The maximal-size PSD constraint involved in (10) under the monomial choice for $v(x)$ is $r\binom{n+k}{k}$.
@@ III-A Consistency Set
Define $h^{0}(A,B)$ as the following residual:
$h_{t}^{0}(A,B)$ $=\hat{x}_{t+1}-A\hat{x}_{t}-B\hat{u}_{t}$ $\forall t\in 1..T-1.$ (11)
The quantity $h^{0}_{t}(A_{\star},B_{\star})$ will be equal to zero at all $t\in 1..T-1$ if there is no noise present in the system. The joint consistency set $\bar{\mathcal{P}}(A,B,\Delta x,\Delta u)\subset\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times m}\times\mathbb{R}^{nT}\times\mathbb{R}^{m(T-1)}$ of plants and noise values consistent with the dynamics (1a) and noise properties (2) is
@@
in which the $(A,B)$ dependence of $h^{0}_{t}$ is implied in notation. The formulation (12) involves bilinear terms $A\Delta x$ and $B\Delta u$, which could lead to nonconvex or possibly disconnected sets $\bar{\mathcal{P}}$ [24]. By defining a matrix $\Xi_{A}\in\mathbb{R}^{n(T-1)\times nT}$ with
$\Xi$ $=[(I_{T-1}\otimes A),\mathbf{0}_{n(T-1)\times n}]+[\mathbf{0}_{n(T-1)\times n},-I_{n(T-1)}],$
and defining $n_{\ell}$ to be the column dimension of $F_{\ell}$ and $G_{\ell}$ for each $\ell$ ($F_{\ell}\in\mathbb{R}^{n_{\ell}\times nT},\ G_{\ell}\in\mathbb{R}^{n_{\ell}\times m(T-1)}$), the consistency set in (12) can be equivalently expressed as
$\bar{\mathcal{P}}:=\ \begin{Bmatrix}[l]\Xi\Delta x+(I_{T-1}\otimes B)\Delta u+h^{0}=0\\ (F_{\ell}\Delta x+G_{\ell}\Delta u,1)\in\mathbb{L}^{n_{\ell}}&\forall\ell\in 1..L\\ \end{Bmatrix}.$ (13)
@@ III-B Full Linear Program
[…]
@@
$\qquad M_{ij}(A,B)-(W\left(A+BK\right)W^{+})_{ij}\geq 0$
$\qquad M_{ij}(A,B)+(W\left(A+BK\right)W^{+})_{ij}\geq 0$
@@
The $W$-superstabilizing linear system in (7) is fulfilled for each $(A,B)\in\mathcal{P}$, as expressed by the $\forall(A,B,\Delta x,\Delta u)\in\bar{\mathcal{P}}$ quantification in (15a). ∎
Remark 1
Under the assumption that $\bar{\mathcal{P}}$ is compact, the function $M$ can be selected to be continuous (Lemma 3.1 of [25]) and even polynomial (Lemma 3.2 of [25]). The same type of linear system structure for superstabilization in (15b)-(15d) is used (with $W=I_{n}$) as in [25], but the set $\bar{\mathcal{P}}$ in [25] is defined by polytopic-bounded noise rather than quadratically-bounded noise (12).